40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Feb 7 15:46:42 2013, Jenne, Update, Locking, PR2 flipped 
    Reply  Fri Feb 8 02:12:14 2013, Jenne, Update, Locking, PRMI work 
       Reply  Fri Feb 8 02:38:04 2013, Koji, Update, Locking, PRMI work 
       Reply  Fri Feb 8 11:01:18 2013, Jamie, Update, Locking, PRMI work 
          Reply  Fri Feb 8 12:39:32 2013, yuta, Update, Locking, PRMI work 
          Reply  Fri Feb 8 15:53:48 2013, Jenne, Update, Locking, PRMI work 
             Reply  Fri Feb 8 17:41:34 2013, Jenne, Update, Locking, PRMI work 
                Reply  Fri Feb 8 19:39:02 2013, Koji, Update, Locking, PRMI work 
                Reply  Fri Feb 8 20:05:15 2013, Jenne, Update, Locking, PRMI work 
Message ID: 8024     Entry time: Thu Feb 7 15:46:42 2013     Reply to this: 8030
Author: Jenne 
Type: Update 
Category: Locking 
Subject: PR2 flipped 

More correctly, a different G&H mirror (which we have a phase map for) was put into the PR2 TT, backwards.

Order of operations:

* Retrieve flat test G&H from BS chamber.  Put 4th dog clamp back on BS optic's base.

* Remove flat G&H from the DLC mount, put the original BS that was in that mount back.  Notes:  That BS had been stored in the G&H's clean optic box.  The DLC mount is engraved with the info for that BS, so it makes sense to put it back.  The DLC mount with BS is now back in a clean storage box.

* Remove PR2 TT from ITMX chamber.

* Remove suspension mounting block from TT frame, lay out flat, magnets up, on lint-free cloth on top of foil.

* Remove former PR2 G&H optic.

* Put what was the flat G&H test optic into the PR2 optic holder, with AR surface at the front.

* Put PR2 suspension block back onto TT frame.

* Put PR2 assembly back in the chamber, solidly against the placement reference blocks that Evan put in last Thursday.

* Close up, clean up, put labels on all the boxes so we know what optic is where.

 

Why the switcho-changeo?  We have a phase map for the G&H that is the new PR2, and a measured RoC of -706m, surface rms of 8.7nm.  Now, we can measure the former PR2 and see how it compares to our estimate of the RoC from the cavity measurements we've taken recently.

ELOG V3.1.3-